OpenStudy (anonymous):

10. The trig functions cosine and secant are related through a(n) _______ trig identity. A. Pythagorean B. reciprocal C. inverse D. quotient

OpenStudy (abdullahm):

\(\sf\Large sec(x)=\frac{1}{cos(x)}\) What do you think it is?

OpenStudy (anonymous):

@AbdullahM can you help me out

OpenStudy (anonymous):

a.

OpenStudy (abdullahm):

No; Phytagorean trig identity is based off of \(\sf a^2+b^2=c^2\) And a Phytagorean Trig identity example is: \(\sf sin^2(x)+cos^2(x)=1\)

OpenStudy (anonymous):

so can you help me @AbdullahM

OpenStudy (anonymous):

im so lost :(

OpenStudy (abdullahm):

What is cos to sec? Is it the inverse? reciprocal? quotient?

OpenStudy (anonymous):

idk

OpenStudy (abdullahm):

If anything else, you can seach up inverse trig identity reciprocal trig identity and see if \(\sf\Large sec(x)=\frac{1}{cos(x)}\) comes under it.

OpenStudy (anonymous):

CAN YOU HELP ME @AbdullahM !!!!!!!!!!!!!!!!!!!!!!!!!!!!!1

OpenStudy (abdullahm):

Tag me in your post please.

OpenStudy (anonymous):

who are you talking to

OpenStudy (abdullahm):

you

OpenStudy (anonymous):

the question is the same as @coolswag and the answers are the same

OpenStudy (abdullahm):

ok

OpenStudy (anonymous):

so can you help me

OpenStudy (abdullahm):

I'll try.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

so......................................................................................................................................... did you find the answer @AbdullahM

OpenStudy (abdullahm):

Please go to the other post where I'm helping :)